OpenStudy (anonymous):

Choose the point which is a solution to the system of inequalities:6x - 5y > -35 3x + y 2 Answers: (4, 7) (-2, 3) (6, -5) (-9, 2)

OpenStudy (anonymous):

can you write the inequalities again please ? and tell me what you got first

OpenStudy (anonymous):

Y less than with a line underneath 2

OpenStudy (anonymous):

My answer is B

OpenStudy (anonymous):

so 6x-5y > -35 3x+y <= 2

OpenStudy (anonymous):

I'm wrong:(

OpenStudy (anonymous):

i wrote it correctly ? this is the question ?

OpenStudy (anonymous):

I worked through it on my own and got the wrong answer I mean

OpenStudy (anonymous):

ok but the question is ? : 6x-5y > -35 3x+y <= 2

OpenStudy (anonymous):

Yes love

OpenStudy (anonymous):

ok how did you try to do it ? graphed it ?

OpenStudy (anonymous):

or you tried to plug in values from the solutions

OpenStudy (anonymous):

The latter

OpenStudy (anonymous):

I know I suck

OpenStudy (anonymous):

in fact i think B is the correct one as well

OpenStudy (anonymous):

:0000

OpenStudy (anonymous):

A .(4, 7) 6*4-5*7 > -35 ? 24 - 35 > -35 ? -11 > -35 good! 3*4 + 7 <= 2 ? no! so A is not the answer so we check C (6, -5) 6*6 - 5*(-5) > -35 ? 36 + 25 > -35 good! 3*6 - 5 <= 2 no! so C is not the answer so we check D : (-9, 2) 6*-9 -5*2 >-35 -54 -10 >-35 no! D is not the answer B: (-2, 3) 6*-2 -5*3 > -35 -27 > -35 good! 3*-2 + 3 <=2 -6 + 3 <=2 -3 <=2 correct ! B is the answer

OpenStudy (anonymous):

ill be back in 20 min

OpenStudy (anonymous):

im back but you are not here anymore .. tell me if you understood it and why you said that it is not B ?